A room is 6m long, 4.5m wide and 4m high. The 4 walls and ceilings are to be painted. If the doors and windows have an area of 14m^2, what is the area to be painted?
My answer is 111m^2 but I didn't know if its correct.

Found 2 solutions by rfer, jim_thompson5910:
Answer by rfer(12731)

2*6*4+2*4.5*4+6*4.5-14=
48+36+27-14=97 sq m
Don't for get to remove the window and door area.

Answer by jim_thompson5910(28717)

Area of the ceiling: 6*4.5 = 27 square meters

Area of longer wall: 6*4=24 square meters

There are 2 longer walls. So the total area of the longer walls is 24+24=48 square meters.


Area of shorter wall: 4.5*4 = 18 square meters

Again, there are 2 shorter walls. So the total area of the shorter walls is 2*18=36 square meters.


Add up the individual areas to get: 27+48+36=111


So the total area of the ceilings and the 4 walls is 111 square meters.

Now we need to subtract off 14 square meters (since we're NOT going to paint over the windows or the doors) to get 111-14 = 97


So the total area to be painted is 97 square meters.


You did great. You just needed to do that last step.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website.

Jim